OpenStudy (anonymous):

what is a simpler form of the expression mutiply (2n+2)(2n-2)

OpenStudy (anonymous):

This is the difference of two squares which is x^2-y^2 =(x+y)(x-y). So...4n^2-4

OpenStudy (anonymous):

4n^2-4

OpenStudy (anonymous):

because (a+b)(a-b)=a^2-b^2 in this equation, your a=2n and b=2 (2n)^2=2*2*n*n=4n^2 (2)^2=2*2=4

OpenStudy (anonymous):

And I wouldn't call it a simplier form but rather just state "another" form but if that's what the question asked... then :).